OpenStudy (idealist10):

Solve the initial-value problem y'+(3/x)y=(3x^4*y^2+10x^2*y+6)/(x^3(2x^2*y+5)), y(1)=1.

OpenStudy (idealist10):

@SithsAndGiggles

OpenStudy (idealist10):

\[y=\frac{ 3x^4 y^2+10x^2 y+6 }{ x^3 (2x^2 y+5) }-\frac{ 3y }{ x }\]

OpenStudy (idealist10):

Wait, supposed to be y' instead of y ^^

OpenStudy (idealist10):

\[y'=\frac{ 6-3x^4 y^2-5x^2 y }{ x^3 (2x^2 y+5) }\]

OpenStudy (anonymous):

Slight mistake with the above. Missing a factor of \(x\). This equation looks a lot like the one we did a while back, so I suspect it'll use the same process. \[\begin{align*} y'&=\frac{ 6-3x^4 y^2-5x^2 y }{ x^3 (2x^2 y+5) } \end{align*}\] Letting \(\color{red}{v=yx}\), you have \(y'=\dfrac{xv'-v}{x^2}\), and the eq. is \[\begin{align*} \frac{xv'-v}{x^2}&=\frac{ 6-3v^2-5v }{ x^3 (2v+5) }\\\\ xv'-v&=\frac{ 6-3v^2-5v }{ x(2v+5) }\\\\ xv'&=\frac{ 6-3v^2-5v +vx(2v+5)}{ x (2v+5) }\\\\ v'&=\frac{ 6-3v^2-5v +vx(2v+5)}{ x^2 (2v+5) }\end{align*}\]

OpenStudy (idealist10):

Didn't we have the same y'? \[y=\frac{ 6-3x^4 y^2-5x^2 y }{ x^3 (2x^2 y+5) }\]

OpenStudy (anonymous):

Yes, I was referring to the post I deleted. Sorry for the confusion.

OpenStudy (idealist10):

So if v=yx, then why do you have \[y=\frac{ 6-3v^2-5v }{ x^3 (2v+5) }\]

OpenStudy (idealist10):

Shouldn't v=(x^2)(y)?

OpenStudy (anonymous):

Yeah you're right! This is what happens when I don't work out a problem like this on paper. Errors everywhere. Sorry again!

OpenStudy (anonymous):

So yeah, \[v=yx^2~~\implies~~y'=\frac{x^2v'-2xv}{x^4}\] So the equation should be \[\begin{align*} \frac{x^2v'-2xv}{x^4}&=\frac{ 6-3v^2-5v }{ x^3 (2v+5) }\\\\ x^2v'-2xv&=\frac{x( 6-3v^2-5v )}{2v+5 }\\\\ xv'-2v&=\frac{ 6-3v^2-5v }{2v+5 }\\\\ xv'&=\frac{ 6-3v^2-5v +2v(2v+5)}{2v+5 }\\\\ \frac{2v+5}{6+5v+v^2}~dv&=\frac{dx}{x} \end{align*}\] a relatively simple separable equation.

OpenStudy (idealist10):

Thank you!